The Sum of two numbers is 20. Three times the
Smaller is equal to two times the larger. Find the two numbers.

Answer by mananth(16949)

let x be one number
other number be y
x+y=20
3x=2y
3x-2y=0
..
Add the two equations
2x+2y+3x-2y=40
5x =40
x=8
..
so y=12